Solve the following system of equations using the substitution method:
4x - y = -6
x + 2y = 3
please show work

Answer by elima(1433)

4x - y = -6
x + 2y = 3
First solve one of the equations for either x or y, we will go ahead and solve the second one for x;
x +2y =3
x = 3-2y
now substitute the this equation for x in the first equation;
4(3-2y)-y=-6
12-8y-y=-6
12-9y=-6
-9y = -6-12
-9y = -18
y = 2
Now that we have what y equals lets plug that in for y in the second equation;
x + 2(2) = 3
x + 4 = 3
x = 3-4
x = -1
So we have;
(-1,2)
:)